1 Use the discriminant to determine whether the quadratic equation has two unequal real solutions, a repeated real solution, or no real solution without solving the equation.
5x2 - 2x - 1 = 0

repeated real solution
two unequal real solutions
no real solution
2 Write the general form of the equation of the circle with radius r and center (h, k).
r = 2; (h, k) = (4, -4)

x2 + y2 + 8x - 8y + 20 = 0
x2 + y2 + 8x + 8y + 20 = 0
x2 + y2 - 8x - 8y + 20 = 0
x2 + y2 - 8x + 8y + 20 = 0

3 Write the standard form of the equation of the circle with radius r and center (h, k).
r = 10; (h, k) = (4, -10)

(x + 4)2 + (y - 10)2 = 100
(x + 4)2 + (y - 10)2 = 10
(x - 4)2 + (y + 10)2 = 100
(x - 4)2 + (y + 10)2 = 10


Found 2 solutions by user_dude2008, stanbon:
Answer by user_dude2008(1862)

1
Answer: two unequal real solutions

2

(x-4)^2+(y+4)^2=4
x^2-8x+16+y^2+8y+16-4=0
x^2+y^2-8x+8y+28=0


3
Answer: (x - 4)2 + (y + 10)2 = 100

Answer by stanbon(75887)

1 Use the discriminant to determine whether the quadratic equation has two unequal real solutions, a repeated real solution, or no real solution without solving the equation.
5x2 - 2x - 1 = 0
---
b^2-4ac = 2^2-4*5*-1 > 0, so there are 2 unequal Real Number solutions.
repeated real solution
two unequal real solutions
no real solution
----------------------------------

2 Write the general form of the equation of the circle with radius r and center (h, k).
r = 2; (h, k) = (4, -4)
----
(x-4)^2 + (y+4) = 2^2
x^2-8x+16 + y^2+8y+16 = 4
x^2 + y^2 -8x+8y + 28 = 0
==========================
x2 + y2 + 8x - 8y + 20 = 0
x2 + y2 + 8x + 8y + 20 = 0
x2 + y2 - 8x - 8y + 20 = 0
x2 + y2 - 8x + 8y + 20 = 0
===================================
3 Write the standard form of the equation of the circle with radius r and center (h, k).
r = 10; (h, k) = (4, -10)
(x-4)^2 + (y+10)^2 = 100
===============================
(x + 4)2 + (y - 10)2 = 100
(x + 4)2 + (y - 10)2 = 10
(x - 4)2 + (y + 10)2 = 100
(x - 4)2 + (y + 10)2 = 10
===============================
Cheers,
Stan H.
===========